Assume that Austin Water purchases surface water from the Lower Colorado River Authority at a cost of $120,000 per month in the months of February through September. Instead of paying monthly, the utility makes a single payment of $880,000 at the end of the year (i.e., end of December) for the water it used. The delayed payment essentially represents a subsidy by the Authority to the water utility. At an interest rate of 3% per year compounded monthly, what is the amount of the subsidy?


1
Expert's answer
2021-06-03T18:36:21-0400

At an interest rate of 3% per year compounded monthly, using the annuity formula the amount of the subsidy is:

"S = P * ((1 + i)^n \u2013 1)\/i - 880,000 = 120,000\u00d7((1 + 0.03\/12)^8 - 1)\/(0.03\/12) - 880,000 = 88,442.13."